OpenStudy (kj4uts):

Suppose that a population parameter is 0.3 and many large samples are taken from the population. If the sample proportions are normally distributed, with 99.7% of the sample proportions falling between 0.018 and 0.582, what is the standard deviation of the sample proportions? A. 0.094 B. 0.074 C. 0.064 D. 0.084

OpenStudy (kj4uts):

@perl I am following the steps that you gave me last week for a question like this but how do I find the z score.

OpenStudy (perl):

$$ \Large \rm { Margin ~of ~ Error = Z_c \cdot \sqrt{\frac{p(1-p)}{n}} }$$

OpenStudy (perl):

the margin of error is half of the given interval

OpenStudy (kj4uts):

is p the 0.3?

OpenStudy (perl):

yes

OpenStudy (kj4uts):

what is n then

OpenStudy (kj4uts):

.997?

OpenStudy (perl):

n is not given

OpenStudy (perl):

n is the sample size

OpenStudy (perl):

but you know margin of error, and you know Zc

OpenStudy (kj4uts):

so half of 0.3 is 0.15

OpenStudy (perl):

$$ \Large \rm { Margin ~of ~ Error = Z_c \cdot \sqrt{\frac{p(1-p)}{n}} \\~\\Margin ~of ~ Error = \frac{0.582 - 0.018}{2} \\ Z_{99.7} = 3 } $$

OpenStudy (kj4uts):

0.282

OpenStudy (perl):

so it turns out you dont need to know n, since the right hand side is the standard deviation

OpenStudy (kj4uts):

A. 0.094 so all you really have to do is subtract the sample proportions divide by 2 to get 0.282 and divide over 3. Where do you get the 0.7?

OpenStudy (perl):

$$\Large \rm { Margin ~of ~ Error = Z_c \cdot \color{red}{Standard~Deviation} \\Margin ~of ~ Error = Z_c \cdot \color{red}{\sqrt{\frac{p(1-p)}{n}} } \\~\\\large Margin ~of ~ Error = \frac{0.582 - 0.018}{2}= 0.282 \\ Z_{99.7} = 3 \\~\\0.282 = 3 \cdot \color{red}{\sqrt{\frac{0.3\cdot 0.7}{n}}} \\\frac{0.282}{3} = \color{red}{\sqrt{\frac{0.3\cdot 0.7}{n}}} } $$

OpenStudy (perl):

I put the standard deviation in red

OpenStudy (perl):

if p = .3 then 1- p = .7

OpenStudy (kj4uts):

Ok I see

OpenStudy (kj4uts):

I got 0.094 as the answer thank you for explaining this to me again I still was not 100% sure how to solve these types of problems. Thank you :)